1.      Given that sin α =  and sin β = , find 𝑡𝑎𝑛 (α + β) if both α and β are in QIV.


1
Expert's answer
2021-03-04T06:02:51-0500


"\sin \alpha=a, -1\leq a\leq0"

"\sin\beta=b, -1\leq a\leq0"

"\cos \alpha=\sqrt{1-\sin^2 \alpha}=\sqrt{1-a^2}"

"\cos\beta =\sqrt{1-\sin^2 \beta}=\sqrt{1-b^2}"


"\tan(\alpha+\beta)=\dfrac{\tan\alpha+\tan\beta}{1-\tan\alpha\tan\beta}"

"=\dfrac{\dfrac{a}{\sqrt{1-a^2}}+\dfrac{b}{\sqrt{1-b^2}}}{1-\dfrac{a}{\sqrt{1-a^2}}\dfrac{b}{\sqrt{1-b^2}}}"

"=\dfrac{a\sqrt{1-b^2}+b\sqrt{1-a^2}}{\sqrt{1-a^2}\sqrt{1-b^2}-ab}"

"\tan(\alpha+\beta)=\dfrac{a\sqrt{1-b^2}+b\sqrt{1-a^2}}{\sqrt{1-a^2}\sqrt{1-b^2}-ab}"